OpenStudy (anonymous):

Simplify 25 – 10 / 5 + 3 x 2

OpenStudy (jamesj):

What is 10/5 and what is 3 x 2 ?

OpenStudy (anonymous):

lol, 2 and 6

OpenStudy (jamesj):

Right ... so 25 - 10/5 + 3x2 = 25 - ....

OpenStudy (anonymous):

i'm still confuseeedddd

OpenStudy (anonymous):

wait so is the answer 29?

OpenStudy (jamesj):

Well, now you know that value of 10/5 and 3x2 So substitute them in to your expression here: 25 - 10/5 + 3x2 = 25 - 2 + 6 Yes?

OpenStudy (jamesj):

Yes 29 is the answer.

OpenStudy (jamesj):

because 25 - 2 + 6 = 29. Voila.